OpenStudy (anonymous):

Find the thirtieth term of the following sequence. -6, -4, -2, 0, ... 52 64 66

OpenStudy (freckles):

is there a common difference? that is do you have the following equality holds: -4-(-6)=-2-(-4)=0-(-2) if so, then this an arithmetic sequence let's call it d let's call a1 the first number in your seqence then the nth term of your sequence is given by: \[a_n=a_1+d(n-1)\] so you want to find a30 so replace the n's with 30 and the d with the number you found above and then you will have a30